OpenStudy (honeymonster):

i need to simplify the following fractions..they are posted in the replies..

OpenStudy (honeymonster):

\[\frac{ 6x + 8x ^{2} }{2x ^{2} }\] \[\frac{ 3x+6x ^{2} }{ 3x }\] \[\frac{ 5x-15x ^{2} }{ 10 }\] \[\frac{ 2x ^{3}+4x ^{2}+6x }{2x}\] \[\frac{ 4x ^{3}+6x ^{2} }{ 8x}\] \[\frac{ x ^{2}-x }{ x-1 }\]

OpenStudy (ghazi):

\[\frac{ 6x+8x^2 }{ 2x}=2 x (3+4x)/2x=(3+4x)\] similarly go for other questions